OpenStudy (anonymous):

Please check my answer c: What are the x- and y-intercepts of the line? 2x – 3y = 12 A. x-intercept –6, y-intercept 4 B. x-intercept –4, y-intercept 6 C. x-intercept 4, y-intercept –6 D. x-intercept 6, y-intercept –4 <<<<<<<

OpenStudy (anonymous):

i got D for my answer

pooja195 (pooja195):

2x – 3y = 12 To find the x intercept you would substitute the y value for 0 2x – 3(0) = 12 x=6 Solve for x For the y intercept you would sub in the x value for 0 2(0) – 3y = 12 y=-4 so yes you are correct :-)